Working together, Edward and Frank can wallpaper an apartment in 12 hrs. It would take Frank 30 hours to do the same job by himself. How long would it take Edward to do it alone?

Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

Frank does 1/30th. of the job in 1 hour
Frank and Ed do 1/12Th job in 1 hour
1/12 - 1/30 = part of job done by ed in 1 hour
5-2 / 60
3/60 = 1/20 th of jon in 1 hour
Ed takes 20 hours to do the job alone

Answer by richwmiller(17219)

12/e+12/30=1
12/e=1-12/30
12/e=18/30
e=20
Edward would take 20 hours alone
check
12/20+12/30=1
6/10=+4/10=1
ed does 6/10 and frank does 4/10 of the work